The length of a rectanglar pool is 6 meters less than twice the width. If the pool's perimeter is 126 meters, what are the dimensions?
Answer by TimothyLamb(4379)

L = 2W - 6
2W = L + 6
P = 2L + 2W = 126
---
2L + (L + 6) = 126
3L = 120
---
Answer:
L = 40
W = 23
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php